6. -12 points Tipler6 2.P.058 A sports car accelerates in third gear from 47.1 km/h to 83.9 km/h in 3.9 s (a) What is the average acceleration of the car? m/s2 (b) If the car maintained this acceleration after reaching 83.9 km/h, how fast would it be moving 2.0 seconds later? km/h eBook
Answer #1

a)

vi = 47.1 km/h

= 47.1*1000 m / 3600 s

= 13.08 m/s

vf = 83.9 km/h

= 83.9*1000 m / 3600 s

= 23.31 m/s

t = 3.9 s

use:

vf = vi + a*t

23.31 = 13.08 + a*3.9

a*3.9 = 10.23

a = 2.62 m/s^2

Answer: 2.62 m/s^2

b)

vi = 83.9 km/h

= 83.9*1000 m / 3600 s

= 23.31 m/s

a = 2.62 m/s^2

t = 2.0 s

use:

vf = vi + a*t

= 23.31 + 2.62*2.0

= 23.31 + 5.24

= 28.6 m/s

= 28.6*3600/1000 km/h

= 103 km/h

Answer: 103 km/h